OpenStudy (anonymous):

is 4x^2+12x+9 a trinomial square?

OpenStudy (anonymous):

if it is it must look like \((a+b)^2\) and since \(3^2=9\) and \((2x)^2=4x^2\) try \((2x+3)^2\) and see if it works

OpenStudy (anonymous):

Its a true or false answer though..

OpenStudy (anonymous):

yes i see that if \((2x+3)^2=4x^2+9x+12\) then it is true. if not, then it is false you have to multiply out and see if it works

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

btw it is true